OpenStudy (anonymous):

I need help with the following: Find the slope of the line that goes through the given points. (-1,-2) and (-2,-1) Select the correct choice below and if necessary, fill in the answer box. a) the slope of the line is _____________ b)the slope is undefined.

OpenStudy (anonymous):

you can use the point slope form

OpenStudy (anonymous):

point slope formula y-y1 = m(x-x1) m = slope

OpenStudy (anonymous):

I am thinking that the slope is undefined??

OpenStudy (anonymous):

your can try it out

OpenStudy (anonymous):

I am lost! :(

OpenStudy (anonymous):

do you know how to plug in the values for y and x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I should...I am just lost & seemed to have forgotten everything

OpenStudy (anonymous):

this is how you use the point slope formula y-y1 = m (x-x1) (x , y) this corresponds to 1 the x1 and y 1 do not get them mixed up (X, Y) this corresponds to the x and y i am not capitalizing purposely only to have you get used to seeing this

OpenStudy (anonymous):

So is the slope 2 ?

OpenStudy (anonymous):

could you show your work?

OpenStudy (anonymous):

no, to be honest I am guessing...seemed to me that the negative numbers cancel out the positive?? Am I correct that the slope is 2?

OpenStudy (anonymous):

just follow the thing i typed up there if it is confusing ill type it a better way :P

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

just be careful around those neg -neg = positive